OpenStudy (happyfeet8040):

Metal will be given :) Solve for x: 2 over 3 (x - 4) = 2x. 2 -2 -8 -4

OpenStudy (mokeira):

\[\frac{ 2 }{3 }(x-4)=2x\] divide both sides by 2/3 \[(x-4)=3x\] collect like terms \[x-3x=4\] \[-2x=4\] \[x=-2\]

OpenStudy (happyfeet8040):

Thank you so much! :)

OpenStudy (mokeira):

glad to help